Observing Log for 2015 Mar 25

00:35 - switch LBC to MODS

01:00 - acquisition of mods.4.z5laeC

01:10 - start science observations. Seeing 0.8".

02:15 - start second sequence of science observations. Applied 0.5" offset along slit. Still clear with 0.8" seeing.

02:30 - getting about 1 mag extinction now from some passing high clouds.03:00 - clearing up, seeing to 0.6"

03:25 - start third sequence. Still occasional passing clouds, but mostly clear with 0.7" seeing.

04:30 - start fourth sequence. Clouds thickening. Counts starting to drop.

04:50 - recovering somewhat, seeing up to 1"

05:20 - cut off last image of sequence due to sky brightness

05:27 - BD+33

-- %USERSIG{Ian McGreer - 2015-03-25}%

Comments


 
Topic revision: r1 - 05 Apr 2015, EiichiEgami
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback